OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

isolate x and y, and remember y = mx+b where m is the slope

OpenStudy (anonymous):

5x - 3y = -6 5x+6=3y y=(5/3)x + 6/3 i.e. y=(5/3)x + 2 comparing it with y=mx+c we find slope m= 5/3 y-intercept = c= 2

OpenStudy (anonymous):

the line intercepts y when x is equal to 0 so plug that in and you will have your answer